OpenStudy (anonymous):

In a precision bombing attack there are 50% chance that any one bomb will strike the target. Two direct hits are required to destroy the target completely. How many bombs must be dropped to give a 99% chance or better of completely destroying the target?

terenzreignz (terenzreignz):

You've got some patience if you've actually waited this long... or have you already gotten it? :)

OpenStudy (anonymous):

true, I am still curious ;P

terenzreignz (terenzreignz):

Well it's simple... after n bombs (talk about overkill) What are the chances that ALL of them fail?

terenzreignz (terenzreignz):

(either overkill or terrible aim)

OpenStudy (anonymous):

that would be (0.5)^n

terenzreignz (terenzreignz):

(imagine the collateral damage)

terenzreignz (terenzreignz):

But that's right :D The chances of all of them failing are (this way is fancier:) \[\Large \frac1{2^n}\]

terenzreignz (terenzreignz):

So... the event is where none of the bombs hit their mark.... The COMPLEMENT, then, of that event, means that at least one of the bombs finds its target.

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

That's what we want... after dropping n-bombs, the chances of at least one bomb hitting its mark would be \[\Large 1- \frac1{2^n}\]

terenzreignz (terenzreignz):

But then again, we need two bombs to hit... hmm

OpenStudy (anonymous):

yeah right :)

terenzreignz (terenzreignz):

err... to completely destroy the target...

OpenStudy (anonymous):

if enough bombs are dropped, we would have 1% for none of them hitting and 99% for at least one of them I'm not really sure how to determine the chance of two of them hitting.......

terenzreignz (terenzreignz):

Yes... let me have a good think...

OpenStudy (anonymous):

well, 2 or more it only asks AT LEAST two, and not precisely two

terenzreignz (terenzreignz):

What are the chances of only one bomb hitting its mark, then?

OpenStudy (anonymous):

good question

terenzreignz (terenzreignz):

that means you haven't gotten round to trying it yet? :D

OpenStudy (anonymous):

no, not at all

OpenStudy (anonymous):

we don't care for the order, if one of them hits out of n it is 1 hit

terenzreignz (terenzreignz):

n-bombs dropped (seriously, what did these people do to deserve that) the chances of ONLY the first bomb hitting its mark is 0.5...times the probability that the n-1 other bombs don't hit their mark right?

OpenStudy (anonymous):

first bomb to hit is 0.5

terenzreignz (terenzreignz):

Essentially, equal to: \[\Large \frac12 \times Pr(\text{rest of the bombs miss})\]

terenzreignz (terenzreignz):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is one possibility :)

terenzreignz (terenzreignz):

Interestingly, that amounts to \[\Large \frac12 \times \frac1{2^{n-1}}= \frac1{2^n}\]

terenzreignz (terenzreignz):

That's the chances for ONLY the first bomb hitting... what about the chances for only the SECOND bomb hitting? By similar reasoning, it should also be \[\Large \frac1{2^n}\] and the same goes for all n-bombs.

terenzreignz (terenzreignz):

Am I right?

OpenStudy (anonymous):

yes, actually, if we want the first bomb to hit and all others not to hit that's just as unlikely as if we wanted to make them all miss. we made just as many requirements with the same probability. and, if we pick any specific bomb and change the requirement from miss to hit, it's also the same amount of requirements, with the same probability.

terenzreignz (terenzreignz):

It's a result of the 50% chance of hitting for each bomb (stupid bombardier) Anyway, the chances, then, that PRECISELY one bomb hits its mark would be the sum total of all the probabilities for each of the particular bombs... all n of them, and nothing would be subtracted, since these events are disjoint (may not both happen; you can't have ONLY the first bomb hitting and ONLY the second bomb hitting at the same time) Then the chances of exactly one bomb hitting is \[\Large \frac{n}{2^n}\] aye?

OpenStudy (anonymous):

"It's a result of the 50% chance of hitting for each bomb (stupid bombardier) hahaha :)

terenzreignz (terenzreignz):

But I'm right, right?

OpenStudy (anonymous):

yes, n x 1/2^n

OpenStudy (anonymous):

describes one bomb hitting

terenzreignz (terenzreignz):

Thus, the chances of LESS THAN TWO bombs hitting would be \[\LARGE \left.\begin{matrix}\frac1{2^n}&+&\frac{n}{2^n}\\\text{no bombs hit}&&\text{exactly one bomb hits}\end{matrix}\right.\]

OpenStudy (anonymous):

agreed

terenzreignz (terenzreignz):

Strictly speaking, minus the chances that both no bombs hit and exactly one bomb hits, but that's nil.

OpenStudy (anonymous):

what is nil ?

terenzreignz (terenzreignz):

zero XD

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

hmm you're right if - not exactly one hits - not none hits = at least two hit!

terenzreignz (terenzreignz):

I'm right? LOL that tends to happen occasionally :D Now then, after n-bombs are dropped, the chances of the target being destroyed (due to not being able to get good help these days) is given by \[\Large 1 - \left.\frac{n+1}{2^n}\right.\]

terenzreignz (terenzreignz):

For the probability to be 99% or more, then we must have \[\Large 1 - \left.\frac{n+1}{2^n}\right.\ge 0.99\] or equivalently \[\Large \frac{n+1}{2^n}\le\frac1{100}\]

terenzreignz (terenzreignz):

Equivalently still, \[\Large 100n + 100 \le 2^n\]

OpenStudy (anonymous):

thanks a lot !!

terenzreignz (terenzreignz):

What's the answer?

OpenStudy (anonymous):

actually how do you solve it when n is on both sides ?

terenzreignz (terenzreignz):

You're actually trying to solve for n? Perish the thought, there is no (nice) way to solve for n. What you could do is try a few reasonable values of n, (remember that n could only take positive integer values! ... unless half-a-bomb is somehow feasible) and check where it starts to become true... (try n = 1 (false) n=2 (false ... etc)

OpenStudy (anonymous):

yeah I was planning on just trying the n ;) just wanted to make sure if there's another way

terenzreignz (terenzreignz):

There isn't :P

terenzreignz (terenzreignz):

are you in High School?

OpenStudy (anonymous):

ok let's see if n=5 \[\Large 100(5) + 100 \le 2^5\]\[\Large 600 \le 32\]need more bombs, if n=15 \[\Large 100(15) + 100 \le 2^{15}\]\[\Large 1600 \le 32768\]probably too many bombs n=10\[\Large 100(10) + 100 \le 2^{10}\]\[\Large 1100\le 1024\]n=11\[\Large 100(11) + 100 \le 2^{11}\]\[\Large 1200\le 2048\]we need to throw 11 bombs and we have the 99% chance of sinking the ship

terenzreignz (terenzreignz):

Very good :D And the precise (more or less) answer is \[\Large n\ge 10.118\]

OpenStudy (anonymous):

@terenzreignz I'm actually at college but I can't do math.. did you get the precise answer by guessing as well??

terenzreignz (terenzreignz):

No, silly, I used technology XD

OpenStudy (anonymous):

so you used some solver or something ? what

terenzreignz (terenzreignz):

wolfram :3

OpenStudy (anonymous):

OK :) thanks again for all the help. can understand problem well now

terenzreignz (terenzreignz):

^_^ If you're going to stick around for a while, then call me TJ :D And lol... this was fun :>

OpenStudy (anonymous):

I will see if I can remember it :)

terenzreignz (terenzreignz):

If you can't, I'm totally ignoring you XD LOL JK Have a nice day ^_^

OpenStudy (anonymous):

same!